Suppose supply and demand for air-conditioners in the little Hawaiian town of Paia (Maui) are given by the following equations:

Qs=P-300

QD=2100-3P

where the price is expressed US$, and quantity represents air conditioning units.

a) Please find the equilibrium price and quantity and draw the supply and demand graph. Do not worry about the scale, just represent the relevant information.

b) Please compute the consumer surplus, producer surplus and total surplus and identify them in your graph.

c) Searching for new sources of revenue, the town council places a $40 per unit tax on sellers of air- conditioners. Please draw a new graph showing how the tax changes the equilibrium quantity and prices facing buyers and sellers. Calculate the new equilibrium quantity and prices and label them on your graph.

d) Please compute the consumer surplus, producer surplus and total surplus with the tax and identify them in your graph. Compare these answers to those in part b.

e) What percentage of the tax is paid by producers, and what percentage is paid by consumers? Please explain how this relates to elasticity.
Answer #1

ANSWER supply. demand Os: Qd P-300 2100-3 P. (a) At equilibrium P 200 - 2400-3p 48- 2400 P= 2400 $600 Price $600 0- a ♡ = P-3pace & (6) • Consumer surplay - and below the It is the Area above the equilibrium demand curve. the equilibrium Price Produc3) 21 оо - - - 34% = 240 со р: 2ча • 3 6fo Ф. Р-9% 2 1 - 34 - 220 чnis equilibrium parce - $610, Quantity a at supply + tax sNew Consumer rusplass sx (400 - 610x (2700) 45 xto x 270 & $12150 New producer surplus - 3x (70-340) x (2700) x 230 x 200 & $